HUBBLE SPACE TELESCOPE DAILY REPORT #1766 PERIOD COVERED: 0000Z (UTC) 11/22/96 - 0000Z (UTC) 11/25/96 Daily Status Report as of 330/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed FOS/RD 6912 (FOS Cycle 6: Dark Monitoring) The Faint Object Spectrograph (Red Side) was used to make = measurements of the internal background. All exposures were internals, = and the spectrographic observations were made in the image mode. The = observations were executed as scheduled, and no anomalies were noted. 1.2 Completed FGS 7210 (Transfer Scan Check) The FGSs were used to make astrometric observations of star = INCA221-16 in order to execute a transfer scan check. These observations = will be used to determine the possible multiplicity of three SMOV target = candidates. The observations were executed as scheduled, and no = anomalies were noted. 1.3 Completed Three Sets of WF/PC-2 6253 (Pure Parallel Archive = Program) The WF/PC-2 was used to execute a pure parallel program taking = ultraviolet observations running concurrently with prime science = observations. The intent is to search for extra-galactic sources with = flux excesses at 3000 A0 and galactic sources with excesses at 1600 A0. T= he observations were completed as planned, and no problems were reported. 1.4 Completed WF/PC-2 6188 (WF/PC-2 Cycle 5 Calibration: Dark) The WF/PC-2 was used to make internal dark calibration images. = The observations were completed as scheduled, and no problems were = reported. 1.5 Completed Six Sets of FOS/RD 6914 (FOS Cycle 6: Location of = Spectra and Wavelength Calibration - Red Side) The Faint Object Spectrograph (Red Side) was used to make both a = TALED and a wavelength calibration image. All observations were made = using internal light sources. This proposal will be executed a total of = 12 time during cycle 6 (i.e., once per month). The observations were = executed as scheduled, and no anomalies were noted. 1.6 Completed FOS/RD 6585 (UV Spectra of Representative Stellar = Populations of Age > 2) The Faint Object Spectrograph (Red Side) was used to make = mid-ultraviolet spectrographic observations of nuclear region of nearby = galaxy NGC-7619. This galaxy has a relatively old stellar population. = The objective is to fill in crucial gaps in our understanding of stellar = populations, and thus aid in the modeling of galactic evolution. The = observations were completed as planned, and no anomalies were reported. 1.7 Completed Three Sets of WF/PC-2 7202 (The WF/PC-2 B Band Parallel = Survey: A Systematic and Synoptic Study of Galaxy Formation and Evolution = - Short Orbit Case) The WF/PC-2 was used to make observations as part of a B-Band = Survey to make synoptic study of galaxy formation and evolution. These = observations were taken in parallel with the prime science observations. = They were completed as planned, and no anomalies were reported. 1.8 Completed GHRS/1 6897 (Cycle 6: Condensed-Regimen Integrated = Monitoring Program for Side One of the GHRS) The High Resolution Spectrograph (Side One) was used to used to = perform a Condensed-Reigman Integrated Monitoring Calibration. Object = BD28D4211 was observed. This proposal combines into one program several = aspects of instrument monitoring that ordinarily execute at the same = frequency of two times per year. A Wavelength calibration image was also = taken. The observations were executed as scheduled, and no anomalies = were noted. 1.9 Completed FOS/RD/GHRS/1 5934 (A Search for UV Absorption Lines in = the Intracluster Medium of Abell Cluster 1030) The Faint Object Spectrograph (Red Side) and the High Resolution = Spectrograph (Side One) were both used to make observations of the QSO = B21028+313 in order to look for ultraviolet absorption lines in a search = for an intracluster medium in Abell Cluster 1030. Recent low energy = observations suggest that such a medium exists. A wavelength calibration = image was also taken. The observations were executed as scheduled, and = no anomalies were noted. 1.10 Completed GHRS/1/2 6665 (Galaxy Rotation Curves at Large Radius = Using Lyman-Alpha Absorption Line) The High Resolution Spectrograph (Both Sides One and Two) was used = to make spectrographic observations of the background QSO PG0923+201 = through the disk of two foreground spiral galaxies in order to obtain the = galaxy's rotation curve by studying the structure of the Lyman-Alpha = lines. The observations were executed as scheduled, and no anomalies = were noted. 1.11 Completed Two Sets of GHRS/2 6898 (CRIMP2: Cycle 6 = Condensed-Regimen Internal Monitoring Program) The High Resolution Spectrograph (Side Two) was used to perform an = internal condensed-regimen integrated monitoring program. This proposal = combines into one program several of the instruments monitoring = observations that ordinarily are executed at the same frequency of twice = per year. A wavelength calibration image was also taken. The = observations were completed as schedule, and no anomalies were reported. 1.12 Completed GHRS/2 6628 (The Wind and Photosphere of the Unique DO = White Dwarf RE/J0503-289) The High Resolution Spectrograph (Side Two) was used to make = spectrographic observations of the unique DO Class, helium rich, white = dwarf star REJ0503-28 in order to study its stellar winds and = photosphere. This white dwarf star also has a planetary nebula = surrounding it. This is the only white dwarf star in which on going mass = loss has been observed with the IUE satellite. The observations were = completed as planned, and no problems were reported. 1.13 Completed FOS/BL 6484 (Absorbing Outflows in Quasars and AGNs) The Faint Object Spectrograph (Blue Side) was used to make high = signal-to-noise ratio spectrographic observations of the active galactic = nuclei of REJ0503-28 in order to study its absorbing outflows and = physical conditions. The observations were executed as scheduled, and no = anomalies were noted. 1.14 Completed WF/PC-2 6802 (Cosmology With the Medium Deep Survey) The WF/PC-2 was used to make observations taken in parallel with = the prime science observations as part of the Medium Deep Sky Survey. = The observations were executed as scheduled, and no problems were = reported. 1.15 Completed GHRS/1/2 6522 (Probing the Extended Envelope of the = Zeta Aur Type Binary 32 Cyg.) The High Resolution Spectrograph (Sides One and Two) both were = used to make spectrographic observations of the Zeta Aug class binary = system HD192909 (aka, 32 Cyg) in order to study its extended envelope. = This binary system consists of a B class dwarf star, and a K class = supergiant that eclipses the dwarf. A wavelength calibration image was = also taken. The observations were executed as scheduled, and no = anomalies were noted. 1.16 Completed FOS/BL 6717 (UV Spectroscopy of the Luminous Cooling = Flow Nebula in A2597) The Faint Object Spectrograph (Blue Side) was to be used to make = ultraviolet observations of the nuclear regions within the luminous, = cooling flow, nebula A2597. Just before this 10 and one-half hour = proposal was to begin, the FOS safed and remained that way for the = duration of the scheduled proposal event. See Section 2.3 for details of = the safing and recovery. 1.17 Completed Two Sets of WF/PC-2 6609 (The WF/PC-2 B-Band Parallel = Survey: A Systematic and Synoptic Study ) The WF/PC-2 was used to make B-Band survey observations in = parallel with the prime science observations. A total of 216 exposures = will be taken over a period of time to make a statistical study of the = formation and evolution of the galaxies. The observations were executed = as scheduled, and no anomalies were noted. 1.18 Completed Two Sets of WF/PC-2 6909 (WF/PC-2 Cycle 6 Calibration: = Earth Flats) The WF/PC-2 was used to make observations of the bright Earth to = obtain streak flats which are used to construct high quality flat fields = for the WF/PC-2 filter set. The observations were executed as scheduled, = and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 15 Successful : 15 Scheduled Reacquisitions : 30 Successful : 30 2.2 FHST Updates: Scheduled : 23 Successful : 23 2.3 Operations Notes: The FOS safed at 329/112618Z, caused by a high voltage spike. The = instrument was recovered successfully by real-time procedures at = 229/2258Z, and the instrument then went to high voltage nominally at = approximately 330/0036Z. HSTARs 5804 and 5805 were written to describe = this event. PACOR-II reported one missing and three corrupted packets in a = WF/PC-2 observation packet. STR #1 was re-dumped prior to the next = record session and the data was successfully recovered. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH